OpenStudy (anonymous):

In △JKL, what is the length of ? 33 66 33 16.5

OpenStudy (anonymous):

you need a picture to go with your problem plz

OpenStudy (anonymous):

Length of which side?

OpenStudy (anonymous):

KL

OpenStudy (anonymous):

Are you sure you don't want JL?

OpenStudy (anonymous):

Im sure. the question said it wants the length of KL

OpenStudy (anonymous):

The reason I asked is because the length of KL is not among your choices.

OpenStudy (anonymous):

thats a bummer:( okay thanks anyways

OpenStudy (anonymous):

But the length of JL IS among the choices!

OpenStudy (anonymous):

what is the length of jl?

OpenStudy (anonymous):

It's 66 because the sin 30 = 1/2 = 33/JL

OpenStudy (anonymous):

okay and i have another question i need help with.! can you help?

OpenStudy (anonymous):

With the same diagram?

OpenStudy (anonymous):

no a different diagram

OpenStudy (anonymous):

There are many ropes keeping a hot air balloon from floating away before a balloon race. One of these ropes is fixed to the ground at a 45° angle. Another is fixed to the ground at a 30° angle. If the hot air balloon is 21 feet off the ground, what is the distance between the ground directly underneath the balloon and the second rope rounded to the nearest hundredth of a foot? Type the numeric answer only in the box below. Numerical Answers Expected!

OpenStudy (anonymous):

Then, just hit "best response" hered then start a new problem, and I'll be glad to help. Too confusing otherwise.

OpenStudy (anonymous):

okay i did